Can you please help me with this word problem: Admission prices to a move are $4 for children and $7 for adults. If 75 tickets were sold for $456, how many of each type were sold?
I tried working out the problem but no luck.
Thank you

Answer by checkley77(12569)

A+C=75
A=75-C
7A+4C=456
7(75-C)+4C=456
525=7C+4C=456
-3C=456-525
-3C=-69
C=-69/-3
C=23 IS THE NUMBER OF CHILDREN'S TICKETS SOLD.
75-23=52 IS THE NUMBER OF ADULT TICKETS SOLD.
PROOF:
7*52+4*23=456
364+92=456
456=456